OpenStudy (anonymous):

Add and simplify:

OpenStudy (anonymous):

\[2x+8 \over x-4\] + \[x-7 \over x-4\]

OpenStudy (anonymous):

add in the numerator, keep the denominator the same

OpenStudy (chaise):

When the denominators are the same you can close to ignore them for the addition process of the fractions and put them back later. (2x+8) + (x-7) We can gather like terms now. (2x+8) + (x-7) = 3x+1 We must now add back the denominator we 'ignored' earlier. \[\frac{3x+1}{x-4}\] There isn't much you can simplify here, due to the fact that you have a plus sign in the between terms in both the denominator and the numerator.

OpenStudy (anonymous):

\[2x+8+x-7=3x+1\]